# Learning Object Compliance via Young’s Modulus from Single Grasps with Camera-Based Tactile Sensors

### v1

@@ Abstract
Compliance is a useful parametrization of tactile information that humans often utilize in manipulation tasks. It can be used to inform low-level contact-rich actions or characterize objects at a high-level. In robotic manipulation, existing approaches to estimate compliance have struggled to generalize across object shape and material. Using camera-based tactile sensors, we present a novel approach to parametrize compliance through Young’s modulus $E$. We evaluate our method over a novel dataset of 285 common objects, including a wide array of shapes and materials with Young’s moduli ranging from 5.0 kPa to 250 GPa. Data is collected over automated parallel grasps of each object. Combining analytical and data-driven approaches, we develop a hybrid system using a multi-tower neural network to analyze a sequence of tactile images from grasping. This system is shown to estimate the Young’s modulus of unseen objects within an order of magnitude at 74.2% accuracy across our dataset. This is a drastic improvement over a purely analytical baseline, which exhibits only 28.9% accuracy. Importantly, this estimation system performs irrespective of object geometry and demonstrates robustness across object materials. Thus, it could be applied in a general robotic manipulation setting to characterize unknown objects and inform decision-making, for instance to sort produce by ripeness.
@@ 1 Introduction
Humans use tactile information to characterize and interact with objects [1]. Through touch, we sense dynamic properties such as friction and compliance [2]. In the case of compliance, we detect discrepancies between soft and hard objects through fingertip pressure and deformation [3].
Compliance has been established as a useful parameter in robotics to classify objects and inform safe grasping policies [4]. It can be directly utilized for a wide variety of biomedical applications [5, 6, 7] or sorting produce by ripeness [8, 9, 10]. Compliance has also been shown to be useful for manipulation through haptic teleoperation systems [11]. Moreover, a reliable measure of compliance could greatly benefit an intelligent, LLM-backed planning architecture [12, 13].
Tactile sensors deployed on robotic end-effectors offer a similar capability to detect dynamic properties as human fingers [14]. High-resolution, camera-based tactile sensors like GelSight have demonstrated strong performance in perceiving properties such as friction and compliance [15, 16]. Incorporating these sensors into robotic systems has proven to be beneficial for executing dynamic tasks [17, 18], including manipulating deformable objects such as cloth or cables [19, 20].
@@
Even with the help of high-resolution tactile sensing, compliance is particularly difficult to discern in unstructured environments. An object’s deformation in response to loading depends both on the object’s shape and its compliance. Without known geometry and pose tracking, analytical methods cannot be utilized to rigorously determine compliance. Instead, learning-based implementations have been developed to estimate compliance independent of shape [16, 6, 12, 13]. These methods demonstrate relative success in estimating compliance across geometries. However, there remains a challenge of universally describing compliance for all material types.
@@
To better generalize across both shapes and materials, we present a novel compliance estimation system which incorporates analytical modeling alongside learning. Specifically, we use a neural network to combine CNN encoded tactile images with estimates produced by analytical contact models. This hybrid approach allows us to better generalize across materials given an analytical backbone while using learning techniques to account for complex geometries. We validate our approach with a novel dataset of 285 common objects which includes a wide variety of shapes and materials. Our estimation system is shown to perform well across this large, diverse dataset.
Young’s modulus $E$ is a property that defines a material’s resistance to deformation in the elastic regime [21]. We choose to parametrize compliance using Young’s modulus because it is universally defined for all materials and is commonly utilized in well-established mechanical models. For these reasons, it is a natural choice for quantifying compliance in robotic manipulation.
Given an increased ability to generalize across object shape and material type, our compliance estimation architecture could be applied in general robotic manipulation scenarios to characterize unknown objects. More acutely, it could be applied to tasks such as produce sorting to identify ripeness of fruits and vegetables [13, 8, 10].
@@ 2.1 GelSight Sensors & Contact Modeling
Camera-based tactile sensors, such as GelSight sensors, are designed to track contact geometry in high-resolution using a camera placed behind an elastomer gel pad [15, 22, 23, 24, 25, 26]. They are highly effective in texture recognition [27, 28, 29], shape and pose estimation [30, 31], and measuring contact force [15] and friction [32]. However, reliably modeling contact with these soft sensors remains challenging. Simulators have been developed to generate tactile images and deformed geometry of sensors using pre-computed finite element method (FEM) solutions [33]. Finite element analysis has proven a viable way to model soft contact on hardware and used to estimate forces from tactile data [34, 35]. Though modeling with FEM is accurate, it requires a precise map of object geometry and solutions must be pre-computed offline.
Alternative approaches have been developed using Hertzian mechanics to model soft contact in robotic grasping [36, 37]. This closed-form method offers improved computational efficiency over FEM and simulated methods. Still, it requires assumptions on contact geometry, namely that geometry should be axisymmetric. Hertzian contact models have been utilized for force estimation [38, 6] and surface curvature detection from tactile data [39]. Since they are closed-form, these constitutive models can more easily be inverted to determine object compliance.
@@ 2.2 Compliance Estimation
Estimating the compliance of a contacted object remains a complex problem in robotic manipulation, particularly without assumptions on object shape or material type. Analytical approaches to compliance estimation often lack robustness to variations in geometry. A traditional approach to model compliance is to apply Hooke’s Law using measurements of grasping force and positions [4]. Similar models have been foundational in developing low-level grasping control policies [40, 41]. This method can provide an effective stiffness of grasping. However, it requires assumptions on object shapes and may struggle when handling complex geometries without a known map of each grasped object. By incorporating tactile sensing, it is possible to create higher fidelity compliance estimation models [42, 43, 44]. These methods have a stronger ability to sense and generalize across contact geometry but still struggle to account for complex contact geometries.
Recent learning-based approaches to compliance estimation have demonstrated the ability to generalize across geometry but struggle across a wide-range of material types. Work has been done to create a tactile-informed model to estimate the Shore 00 hardness of an object in contact with a GelSight sensor. The architecture predicts hardness directly from tactile images without any mechanical modeling [45, 16]. Shore 00 hardness is a measure designed to quantify the compliance of soft rubbers [46]. Thus, this method is trained using only rubber objects. It is fundamentally restricted to predicting on this scale and cannot express differences in compliance between more rigid objects. LLM-backed methods have been developed to classify tactile images with binary semantic descriptions of ‘soft’ or ‘hard’ [12, 13]. These large-data methods are promising but still have not been used to assess compliance on a continuous scale across material types. Other data-driven implementations have been developed using tactile sensing to directly measure the ripeness of produce [8, 9, 10]. As before, these methods are restricted by their dataset and labeling method. In all cases, labeling choices prevent models from incorporating analytical modeling alongside learning.
@@ 3.1 Overview
We develop a hybrid system fusing analytical models alongside learning to estimate the Young’s modulus of general unknown objects from a single parallel grasp. Through analytical modeling, we can create a well-founded, preliminary estimate for Young’s modulus. By incorporating learning, we can compensate for assumptions taken by analytical models to better generalize across contact geometry. A large, diverse dataset of 285 common objects is collected to train and evaluate our method. Our approach of combining analytical models with data-driven methods is partially inspired by residual physics architectures [47] and physics-informed tactile models [48].
Our hybrid estimation architecture is displayed in Fig. 2. Tactile images and measurements of force $F$ and gripper width $w$ are fed into Hertzian and elastic analytical models to fit respective estimates of Young’s modulus $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$. In parallel, a selected number $N_{\text{frames}}$ of tactile images are fed into a data-driven model. Features extracted from images are concatenated with grasp measurements into a large fully-connected decoder. This multi-tower architecture is similar to other grasp-based learning architectures [49]. Output features from our learned model are sent to a smaller decoder alongside analytical estimates to produce a final hybrid estimate of Young’s modulus $\hat{E}$.
@@ 3.2 Analytical Contact Models
Simple Elasticity Parallel grasping can be modeled as uniaxial loading using an adaptation of Hooke’s Law. A diagram of parallel grasping is provided in Fig. 3. At every instance $t$, force $F$ is applied to the grasped object, with gripper at width $w$. Traditional methods have used these measurements alone to estimate the compliance or stiffness of grasped objects [4].
With camera-based tactile sensing, we can increase the fidelity of the model by measuring contact area $A$ and depth $d$ [15]. Contact area $A$ is determined by masking depth images with constant threshold of 0.1mm. This threshold was chosen empirically to cancel out potential noise. Without tactile sensing, we could not measure depth ($d(t)=0$) and would need to assume constant contact area ($A(t)=A_{\text{sensor}}$). We use measurements to calculate contact stress $\sigma$ and object strain $\epsilon$, following Eq. 1. We define $t=0$ to be the moment of first contact, where $F(t=0)$ must be greater than an empirically-defined threshold.
@@
Despite its simplicity, this method does not properly consider the geometry of unknown objects. It implicitly assumes objects have a consistent profile along the grasping axis. Furthermore, it does not take advantage of precise information about the deformation of contact geometry that is apparent through high-resolution tactile data.
@@
Hertzian Contact Hertzian contact theory examines the deformation of elastic bodies through contact [50]. In a simple case, contact is assumed to be without friction or adhesion [51, 52]. Unlike the previously-described method, Hertzian mechanics more acutely analyzes geometry through contact. It has been used to model soft contact in robotics [36, 39], including GelSight-like tactile sensors [23, 6]. Precise depth sensing allows us to track surface deformation and apply this model.
GelSight sensors are made with highly elastic silicone rubber [22]. Unless the grasped object is fully rigid, we expect both bodies to deform through contact. This implies the maximum measured penetration depth $d$ will be less than the relative displacement between the finger and object $d<\frac{\Delta w}{2}$. This phenomenon is depicted in Fig. 3. An aggregate modulus $E^{*}$ describes the bodies’ combined resistance to deformation in Eq. 3, which depends on the Poisson’s ratio $\nu$ of each body.
$\frac{1}{E^{*}}=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}+\frac{1-\nu_{\text{obj}}^{2}}{E_{\text{obj}}}$ (3)
Hertzian solutions have been derived for a wide range of contact geometries [53]. Since we are analyzing contact with an unknown object whose shape is not precisely known, we model all contact geometry as spherical. This is a reasonable simplification for convex objects, but may produce bounded errors in estimates over our geometrically diverse dataset. Our neural network will learn to compensate for this intrinsic and repeatable geometric modeling error.
Since we assume contact profiles are spherical, we must calculate contact radius $a$. This is the radius needed for a circle’s contact area to match the observed contact area $A$. Additionally, we can approximate the apparent radius of the unknown object $R$ through observed contact area and gripper displacement [53]. Using these parameters and the method of dimensionality reduction (MDR), we can calculate the sensor’s surface deformation in response to elastic loading [53, 54]. This relates force $F$ and depth $d$ by $E^{*}$. This relation shows that a harder sphere will plunge deeper into the contact plane, assuming constant force. More detailed derivation for this expression and discussion about the MDR is provided in Appendix A.
$a(t)=\sqrt{\frac{A(t)}{\pi}},\quad R(t)=\frac{a^{2}(t)}{\Delta w(t)},\quad d(t)=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}{\left(\frac{3{E^{*}}^{2}F(t)}{32R^{2}(t)}\right)}^{\frac{1}{3}}a(t)$ (4)
With measured data over the duration of a grasp, we apply least-squares to retrieve a best fit value $\hat{E}^{*}$. Then, we use $\hat{E}^{*}$ to compute an estimated modulus for our unknown object $\hat{E}_{\text{hertz}}$, given known properties of our sensor $E_{\text{sensor}}=0.275$ MPa, $\nu_{\text{sensor}}=0.48$. We assume constant $\nu_{\text{obj}}=0.4$ based on the Poisson’s ratio of common materials [55]. After analytical estimates are computed, linear scaling is added to account for potential constant errors across the dataset.
@@ 4 Physical Dataset
To evaluate how well estimation algorithms could generalize, diversity of shape and material was prioritized in acquiring objects. Thus, we mold 20 custom objects out of silicone to replicate the exact shape of other objects in the dataset with vastly different material composition. Molded objects are displayed in Fig. 4b. Training on data collected across object replicas of different materials will help prevent overfitting to shapes and textures. Moreover, there is incentive to bias towards rubber objects because it is clear from Eq. 5 that it will be easiest to discriminate between Young’s moduli nearest to our sensor’s. The dataset’s composition is further detailed in Appendix D.
As objects were gathered, they were labeled with their Young’s moduli. For harder objects, like those made of metals and plastics, this was done by referencing an engineering database of material properties [57], as all objects are made of known materials. For softer objects, like those made of rubber, Shore hardness was measured and converted to Young’s modulus. Shore hardness is converted to Young’s modulus using Gent’s model of hardness and other established methods [46, 58]. These models provide a reliable measure of object compliance on a logarithmic scale. The dataset includes materials ranging from foam to steel, with Young’s moduli from 5.0 kPa to 250 GPa. For reference, our sensor’s Young’s modulus is 0.275 MPa.

### v2

# Learning Object Compliance via Young’s Modulus from Single Grasps using Camera-Based Tactile Sensors

@@ Abstract
Compliance is a useful parametrization of tactile information that humans often utilize in manipulation tasks. It can be used to inform low-level contact-rich actions or characterize objects at a high-level. In robotic manipulation, existing approaches to estimate compliance have struggled to generalize across object shape and material. Using camera-based tactile sensors, proprioception, and force measurements, we present a novel approach to estimate object compliance as Young’s modulus $E$ from parallel grasps. We evaluate our method over a novel dataset of 285 common objects, which includes a wide array of shapes and materials with Young’s moduli ranging from 5.0 kPa to 250 GPa. Data is collected over automated parallel grasps of each object. Combining analytical and data-driven approaches, we develop a hybrid system using a multi-tower neural network to analyze a sequence of tactile images from grasping. This system is shown to estimate the Young’s modulus of unseen objects within an order of magnitude at 74.2% accuracy across our dataset. This is a drastic improvement over purely analytical and data-driven baselines which exhibit 28.9% and 65.0% accuracy respectively. Importantly, this estimation system performs irrespective of object geometry and demonstrates robustness across object materials. Thus, it could be applied in a general robotic manipulation setting to characterize unknown objects and inform decision-making, for instance to sort produce by ripeness.
@@ 1 Introduction
General and robust object compliance estimation could significantly enhance robotic perception systems. For instance, a ripe avocado appears visually identical to an unripe avocado, yet their compliance differs greatly. Without the ability to accurately estimate compliance, a robot cannot effectively distinguish between such visually similar objects. Compliance measurement has proven valuable in multiple applications, including identifying the ripeness of fruits and vegetables [1, 2, 3, 4] and in biomedical fields [5, 6, 7]. In the future, a universal compliance estimation architecture could be applied in assembly or recycling processes to inform decisions based on object material.
In robotics, we seek to replicate the human ability to perceive properties like compliance through touch [8]. Humans naturally use tactile information to characterize objects and inform interactions [9]. In the case of compliance, we sense discrepancies between soft and hard objects through fingertip pressure and deformation [10]. High-resolution, camera-based tactile sensors like GelSight are specifically designed to track the deformation of a robotic fingertip during contact [11, 12].
@@
However, even with the help of high-resolution tactile sensing, achieving human-like compliance perception remains challenging, especially in unstructured environments. An object’s deformation in response to loading depends on both its shape and material composition. Thus, without known geometry and pose tracking, analytical methods cannot be utilized to rigorously determine compliance [13]. Alternatively, learning-based implementations have been developed to estimate compliance from tactile images alone without the use of mechanical modeling [14, 5, 15, 1]. While these methods have demonstrated relative success in estimating compliance independent of shape, they still face limitations in describing compliance across a wide range of materials.
@@
We aim to create a compliance estimation system that is robust to both shape and material by integrating analytical modeling with machine learning. Young’s modulus ($E$) is a fundamental property in mechanical modeling that offers a measure of compliance for nearly all materials [16]. By parameterizing compliance using Young’s modulus, we can seamlessly incorporate analytical models and create estimates of compliance on a universal scale. Our hybrid approach leverages the strengths of analytical modeling to better generalize across materials, while using learning techniques to address complexities posed by diverse geometries. Specifically, we use a neural network to fuse CNN-encoded tactile images with estimates from analytical contact models. We train our approach across a novel dataset of 285 common objects, including a wide variety of shapes and materials.
Our primary contribution is a novel compliance estimation archiecture which demonstrates increased accuracy over the current state-of-the-art. Additionally, we contribute a novel dataset which could be used in the development of future compliance estimation architectures. Using our diverse dataset, we validate that our hybrid estimation approach can generalize well across object shape and material. Thus, our estimation system could be applied in general robotic manipulation to characterize unknown objects and inform decision-making. This is especially beneficial in scenarios where visually similar objects may have distinct material composition.
@@ 2.1 GelSight Sensors & Contact Modeling
Camera-based tactile sensors, such as GelSight sensors, are designed to track contact geometry in high-resolution using a camera placed behind an elastomer gel pad [12, 17, 18, 19, 20, 21]. They are highly effective for sensing texture [22, 23, 24], contact force [12], friction [25], and object shape and pose [26, 27]. However, reliably modeling contact with these soft sensors remains challenging. Simulators have been developed to generate tactile images and deformed geometry of sensors using pre-computed finite element method (FEM) solutions [28]. Finite element analysis has proven a viable way to model soft contact on hardware and used to estimate forces from tactile data [29, 30]. Though modeling with FEM is accurate, it requires a precise map of object geometry and solutions often must be pre-computed offline.
Alternative approaches have been developed using Hertzian mechanics to model soft contact in robotic grasping [31, 32]. This closed-form method offers improved computational efficiency over FEM and simulated methods. Still, it requires assumptions on contact geometry, namely that geometry should be axisymmetric. Hertzian contact models have been utilized for force estimation [33, 5] and surface curvature detection from tactile data [34]. These closed-form constitutive models can easily be inverted to determine object compliance.
@@ 2.2 Compliance Estimation
Estimating the compliance of a contacted object remains a complex problem in robotic manipulation, particularly without assumptions on object shape or material type. Analytical approaches to compliance estimation often lack robustness to variations in geometry. A traditional approach to model compliance is to apply Hooke’s Law using measurements of grasping force and positions [13]. Similar models have been foundational in developing low-level grasping control policies [35, 36]. This method can provide an effective stiffness of grasping, but may struggle when applied to unknown objects with complex shapes. By incorporating tactile sensing, it is possible to create higher fidelity compliance estimation models [37, 38, 39]. These methods have a stronger ability to sense and generalize across contact geometry but still struggle to account for complex contact geometries.
Recent learning-based approaches to compliance estimation have demonstrated the ability to better generalize across geometry but struggle when applied over a wide-range of material types. Data-driven architectures have been developed to estimate the Shore 00 hardness of an object in contact with a GelSight sensor directly from tactile images without any mechanical modeling [40, 14]. The Shore 00 hardness scale is designed to quantify the compliance of soft rubbers [41]. This model is trained using exclusively rubber objects and is thus fundamentally limited to this material domain. Alternatively, LLM-backed methods have been developed to classify tactile images with binary semantic descriptions of ‘soft’ or ‘hard’ [15, 1]. These large data models are promising but have not been used to assess finer differences in compliance on a continuous scale. Other data-driven implementations have been developed using tactile sensing to directly measure the ripeness of produce [2, 3, 4]. As before, these methods are restricted by their dataset and labeling method. In all cases, labeling choices inhibit models from integrating analytical modeling alongside learning.
@@ 3.1 Overview
We develop a hybrid system that fuses analytical models alongside learning to estimate the Young’s modulus of general unknown objects from a single parallel grasp. Through analytical modeling, we can create a well-founded, preliminary estimate for Young’s modulus. By incorporating learning, we can compensate for assumptions taken by analytical models to better generalize across contact geometry. A large, diverse dataset of 285 common objects is collected to train and evaluate our method. Our approach of combining analytical models with data-driven methods is partially inspired by residual physics architectures [42] and physics-informed tactile models [43].
Our hybrid estimation architecture is displayed in Fig. 2. After grasping, tactile images and measurements of force $F$ and gripper width $w$ are fed into analytical models to fit respective estimates of Young’s modulus $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$. In parallel, a selected number $N_{\text{frames}}$ of tactile images from a single sensor are processed through a CNN. Features extracted from these images are concatenated with grasp measurements $F$, $w$ into a large fully-connected decoder. This multi-tower network is similar to other grasp-based architectures [44]. Output features from the large decoder are sent to a smaller decoder alongside analytical estimates to produce a final hybrid estimate of Young’s modulus $\hat{E}$.
@@ 3.2 Analytical Contact Models
Simple Elasticity: Parallel grasping can be modeled as uniaxial loading using an adaptation of Hooke’s Law. A diagram of parallel grasping is provided in Fig. 3. At every instance $t$, force $F$ is applied to the grasped object, with gripper at width $w$. Traditional methods have used these measurements alone to estimate the compliance or stiffness of grasped objects [13].
With camera-based tactile sensors, we can increase the fidelity of the model by measuring contact area $A$ and depth $d$ [12]. Contact area $A$ is determined by masking depth images with constant threshold of 0.1mm. This threshold was chosen empirically to cancel out potential noise. Without tactile sensing, we could not measure depth ($d(t)=0$) and would need to assume constant contact area ($A(t)=A_{\text{sensor}}$). We calculate contact stress $\sigma$ and object strain $\epsilon$, following Eq. 1. This model assumes symmetry between parallel fingers. We define $t=0$ to be the moment of first contact, where $F(t=0)$ must be greater than an empirically-defined threshold.
@@
Despite its simplicity, this method does not properly consider object geometry and implicitly assumes a consistent profile along the grasping axis. Furthermore, it overlooks precise information about the deformation of contact geometry that is apparent through high-resolution tactile data.
@@
Hertzian Contact: Hertzian contact theory models the deformation of elastic bodies in contact [45]. In a simple case, contact is assumed to be without friction or adhesion [46, 47]. Unlike the previous method, Hertzian mechanics closely analyzes geometry through contact. It has been applied to model soft contact in robotics [18, 5], including for GelSight-like tactile sensors [31, 34, 18, 5].
GelSight sensors are made with highly elastic silicone rubber [17]. Unless a grasped object is fully rigid, we expect both bodies to deform through contact. This implies the maximum measured penetration depth $d$ will be less than the relative displacement between the finger and object $d<\frac{\Delta w}{2}$. This phenomenon is depicted in Fig. 3. An aggregate modulus $E^{*}$ describes the bodies’ combined resistance to deformation in Eq. 3, which depends on the Poisson’s ratio $\nu$ of each body.
$E^{*}=\left(\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}+\frac{1-\nu_{\text{obj}}^{2}}{E_{\text{obj}}}\right)^{-1}$ (3)
Hertzian solutions have been derived for a wide range of contact geometries [48]. We model the sensor as a flat plane. Although we are analyzing contact with an object whose shape is unknown, we model the geometry of all grasped objects as spherical. This simplification is reasonable for convex objects, but may produce bounded errors over our geometrically diverse dataset. Our neural network can learn to compensate for this intrinsic and repeatable geometric modeling error.
Since we assume contact profiles are spherical, we calculate contact radius $a$ based on observed total contact area $A$. Additionally, we calculate the apparent radius of the unknown object $R$ using contact area and gripper displacement [48]. Using these parameters, we can relate depth $d$ to normal force $F$ via the method of dimensionality reduction (MDR) [48, 49], as formulated in Eq. 4. More detailed derivation for the MDR is provided in Appendix A. With these equations established and measurements over the duration of a grasp, we apply least-squares to retrieve a best fit value $\hat{E}^{*}$.
$a(t)=\sqrt{\frac{A(t)}{\pi}},\quad R(t)=\frac{a^{2}(t)}{\Delta w(t)},\quad\boldsymbol{d}=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}{\left(\frac{3{{\hat{E}}^{*2}}\boldsymbol{F}}{32\boldsymbol{R}^{2}}\right)}^{\frac{1}{3}}\boldsymbol{a}$ (4)
We compute an estimated modulus for our unknown object $\hat{E}_{\text{hertz}}$ through Eq. 5, given known properties of our sensor $E_{\text{sensor}}=0.275$ MPa, $\nu_{\text{sensor}}=0.48$. We assume constant $\nu_{\text{obj}}=0.4$ based on the Poisson’s ratio of common materials [50]. After analytical estimates are computed, linear scaling is added to account for potential constant errors across the dataset.
@@ 4 Physical Dataset
To evaluate how well estimation algorithms could generalize, diversity of shape and material was prioritized in acquiring objects. Thus, we mold 20 custom objects out of silicone to replicate the exact shape of other objects in the dataset with vastly different material composition. Molded objects are displayed in Fig. 4b. Training on data collected across object replicas of different materials will help prevent overfitting to shapes and textures. Moreover, there is incentive to bias towards rubber objects because it is clear from Eq. 5 that it will be easiest to discriminate between Young’s moduli nearest to our sensor’s. The dataset’s composition is further detailed in Appendix E.
As objects were gathered, they were labeled with their Young’s moduli. For harder objects, like those made of plastics and metals, this was done by referencing an engineering database of known material properties [52]. For softer objects, like those made of rubber, Shore hardness was measured and converted to Young’s modulus using Gent’s model of hardness and other established methods [41, 53]. These models provide a reliable measure of object compliance on a logarithmic scale. The dataset includes materials ranging from foam to steel, with Young’s moduli from 5.0 kPa to 250 GPa. For reference, our sensor’s Young’s modulus is 0.275 MPa.